A wire 100cm long and 2mm in diameter has a resistivity of 4.8x10
-8Ωm.
1. Calculate the resistance of the wire?
2. A second wire of the same material has the same weight as the 100m length, but twice its diametric. Evaluate its resistance.
A wire has a resistance of 10.0 ohms at 20.00c and 13.1 ohms at 1000c. Obtain a value for its temperature coefficient of resistance. 

3. What is the potential difference between two points in a circuit if 200J of electrical energy are changed to other forms of energy when 25 coulombs of electric charge pass? If the charge flows in 10 seconds, what is the current? 

A television set shoots out a beam of electrons. The beam current is 10uA. How many electrons strike the TV screen each second? How much charge strike the screen in a minute 

1
Expert's answer
2020-03-02T10:14:17-0500

1) "R=\rho\dfrac{l}{A}" ="4.8*10^{-8}*1/(3.14*10^{-6})=1.52*10^{-2} \Omega"